Two trains 100 km apart each travel towards each other at 10 km per hour. A bumblebee starts flying back and forth between the two trains at 60 km per hour, until the two trains crash into each other. How far does the bumblebee travel?

Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
When each train travels 50km, they will have crashed into each other
Time required for each train to travel 50 km=50/10=5 hr
So the bumblebee gets to fly back and forth for 5 hours before it gets smashed, so:
distance bumblebee travels=60*5=300 km

Hope this helps---ptaylor